OpenStudy (anonymous):

1. Solve for x in the interval [0 , 2 π ). (a) 2sin^2x=sin x (b) 3sin^2x+cos^2x=2 (c) 1+3cosx=cos2x (d) sin2x+sin x=0 (d) sin2x+sin x=0

OpenStudy (anonymous):

what do you not understand here? or you just want copy answer?

OpenStudy (anonymous):

solution (a): \[2\sin ^{2}x-\sin x=0\] \[\sin x(2sinx-1)=0\] \[\implies sinx=0,1/2\] \[\implies x=0,\pi,\pi/6,5\pi/6\]